# The Picard group of the Baily-Borel compactification of the moduli space of polarized K3 surfaces

### v1

@@
Abstract
In this paper, we study the Picard group of the Baily-Borel compactification of orthogonal Shimura varieties. As a result, we determine the Picard group of the Baily-Borel compactification of the moduli space of quasi-polarized K3 surfaces. Interestingly, in contrast to the situation observed in the moduli space of curves, we find that the Picard group of the Baily-Borel compactification is isomorphic to $\mathbb{Z}$.
keywords:
Picard group, moduli space of K3 surfaces, Baily-Borel compactification, vector-valued modular form, Theta series, Weil representation
keywords:
[…]
YEAR
{Frontmatter}
Acknowledgments
Zhiyuan Li is supported by NSFC grant (12121001, 12171090 and 12425105) and Shanghai Pilot Program for Basic Research (No. 21TQ00). Manuel K.-H. Müller acknowledges support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124 as well as support by the LOEWE professorship in Algebra, project number LOEWE/4b//519/05/01.002(0004)/87.
1 Introduction
Exploring the Picard group of moduli spaces stands as a vital and inspiring challenge in moduli theory, rooted in the groundbreaking contributions of Mumford [26] in 1965. Let ${\mathscr{M}}_{g}$ be the moduli space of smooth curves of genus $g$ and ¯$\overline{{\mathscr{M}}}_{g}$ be its Deligne-Mumford compactification. Harer [18] and Arbarello-Cornalba [1] showed that
@@ 1 Introduction
For K3 surfaces, let ${\mathscr{F}}_{g}$ be the (coarse) moduli space of quasi-polarized K3 surfaces of genus $g$. By global Torelli theorem for K3 surfaces, ${\mathscr{F}}_{g}$ is isomorphic to a Shimura variety of orthogonal type. It has been shown in [3, 14] that its Picard group ${\rm Pic}({\mathscr{F}}_{g})\cong\mathbb{Z}^{r_{g}}$ is a free abelian group generated by linear combinations of Noether-Lefschetz divisors. Moreover, its rank is
$r_{g}$ $=\frac{31g+24}{24}-\frac{1}{4}\left(\frac{2g-2}{2g-3}\right)\left(\frac{g}{2}\right)-\frac{1}{6}\left(\frac{g-1}{4g-5}\right)-\frac{1}{6}(-1)^{-\left(\frac{g-1}{3}\right)}$
$-\sum\limits_{k=0}^{g-1}\left\{\frac{k^{2}}{4g-4}\right\}-\sharp\left\{k~{}|~{}\frac{k^{2}}{4g-4}\in\mathbb{Z},0\leq k\leq g-1\right\}$
where $\left(\frac{\cdot}{\cdot}\right)$ is the Jacobi symbol. On ${\mathscr{F}}_{g}$, there exists a canonical Satake compactification, denoted as ¯$\overline{{\mathscr{F}}}_{g}$, as established by the work of Baily and Borel. A natural problem is to figure out the Picard group of ¯$\overline{{\mathscr{F}}}_{g}$. The main result of this paper is
Theorem 1.1 (Theorem 6.5).
${\rm Pic}(\overline{{\mathscr{F}}}_{g})\cong\mathbb{Z}$, which is spanned by an integral multiple of the extended Hodge line bundle $\overline{\lambda}$.
@@
Since $\overline{\lambda}$ is currently only known to be a $\mathbb{Q}$-line bundle, an integral multiple of $\overline{\lambda}$ is needed in the statement of Theorem 1.1. A generator of ${\rm Pic}(\overline{{\mathscr{F}}}_{g})$ can be chosen to be $\overline{\lambda}$ once one can show that $\overline{\lambda}$ is a line bundle.
Our proof relies heavily on the study of the behavior of Heegner divisors on the boundary component of the Baily-Borel compactification established in [9] and a surjectivity result of theta liftings for vector-valued modular forms. In fact, we prove that the Baily-Borel compactification of a Shimura variety of K3 type has Picard number one.
Theorem 1.3 (Corollary 6.4).
Let $M$ be an even lattice of signature $(2,n)$ of K3 type, and let $\Gamma\subset\mathrm{O}(M)(\mathbb{Q})$ be an arithmetic subgroup containing $\widetilde{\mathrm{O}}(M)=\ker(\mathrm{O}(M)\to\mathrm{O}(M^{\vee}/M))$. If $n>10$, then
$\dim_{\mathbb{Q}}{\rm Pic}_{\mathbb{Q}}\left(\overline{{\rm Sh}}^{\mathrm{BB}}_{\Gamma}(M)\right)=1.$
Our result can be applied to the Baily-Borel compactification of moduli spaces of lattice-polarized K3 surfaces, polarized hyper-Kähler varieties, K-stable log Fano pairs and also moduli space of bounded polarized Calabi-Yau pairs (cf. [15, 21, 4]).
@@
We believe that the conditions as stated in Theorem 1.3 can certainly be relaxed, though it might entail more sophisticated developments of our methods. More precisely, we believe that Theorem 1.3 holds for a general even lattice $M$ of signature $(2,n)$ and a possibly more general arithmetic subgroup $\Gamma$, whenever $\dim{\rm Sh}_{\Gamma}(M)\geq 3$.
Another motivation for studying this problem concerns the tautological ring of ${\mathscr{F}}_{g}$. Let
$\mathrm{R}^{\ast}({\mathscr{F}}_{g})\subseteq\mathrm{CH}^{\ast}({\mathscr{F}}_{g})_{\mathbb{Q}}$
be the subring generated by irreducible Noether-Lefschetz cycles. It is conjectured by Oprea and Pandharipande (see [12, Conjecture 2]) that
$\mathrm{R}^{17}({\mathscr{F}}_{g})\cong\mathbb{Q}$
is spanned by $\lambda^{17}$. This is clearly a very challenging problem, especially when $g$ gets larger. As $\mathrm{CH}^{17}({\mathscr{F}}_{g})_{\mathbb{Q}}\cong\mathrm{CH}_{2}({\mathscr{F}}_{g})_{\mathbb{Q}}\cong\mathrm{CH}_{2}(\overline{{\mathscr{F}}}_{g})_{\mathbb{Q}}$, one can regard $\mathrm{R}^{17}({\mathscr{F}}_{g})$ as a subgroup of $\mathrm{CH}_{2}(\overline{{\mathscr{F}}}_{g})_{\mathbb{Q}}$. Then a natural first attempt is to ask whether it is one-dimensional modulo numerical equivalence. Typically, one can consider intersecting the classes in $\mathrm{R}^{17}({\mathscr{F}}_{g})$ with complete intersections in $\mathrm{CH}^{2}(\overline{{\mathscr{F}}}_{g})_{\mathbb{Q}}$ (via cap product in the sense of [17, Chapter 17]), which is well-understood set-theoretically in Shimura geometry. According to the Noether-Lefschetz conjecture, complete intersections of divisors on ${\mathscr{F}}_{g}$ are rationally equivalent to linear combinations of Shimura subvarieties on ${\mathscr{F}}_{g}$, which possess very rich structures. However, Theorem 1.1 indicates that all complete intersections in $\mathrm{CH}^{2}(\overline{{\mathscr{F}}}_{g})_{\mathbb{Q}}$ are proportional to $\overline{\lambda}^{2}$. This first test is now almost trivially passed. Although it suggests supportive evidence for the original conjecture, one still need to find other interesting classes in $\mathrm{CH}^{2}(\overline{{\mathscr{F}}}_{g})_{\mathbb{Q}}$.
1.1 Organization of the paper
In Section 2, we review the basic theory of vector-valued Siegel modular forms including the theta series and Siegel-Eisenstein series. Section 3 and Section 4 contain the most technical part. We use the Bruinier-Stein and Shimura Hecke operators to prove a surjectivity result of theta liftings by adapting the proof of [25] to the odd rank case. Some computations are similar and we give the details in the appendix. Readers with a preference for geometry may choose to skip this part.
In Section 5 and Section 6, we study Heegner divisors on Shimura varieties of orthogonal type and their behavior on the boundary component of the Baily-Borel compactification. The main theorem is proved in Subsection 6.3.
@@ 1.1 Organization of the paper
    $(x,y):=\mathfrak{q}(x+y)-\mathfrak{q}(x)-\mathfrak{q}(y)$
    is a non-degenerate symmetric bilinear form. More generally, for any two elements $\mathbf{u}=(u_{1},...,u_{d}),\mathbf{v}=(v_{1},...,v_{d})\in G^{(d)}$, we define $(\mathbf{u},\mathbf{v}):=\begin{pmatrix}(u_{i},v_{j})_{i,j}\end{pmatrix}\in{\rm Sym}_{d}(\mathbb{Q}/\mathbb{Z})$. The level of $G$ is
    $\ell(G):=\min\{m\in\mathbb{Z}_{>0}\mid~{}m\mathfrak{q}(x)=0\hbox{ for all }x\in G\}.$
    Note that since the bilinear form (,)(-,-)( - , - ) is non-degenerate, for every $\gamma\in G$, we have $\ell(G)\cdot\gamma=0$. Finally, the $p$-rank of $G$ is ${\rm rank}_{p}(G):=\dim_{\mathbb{F}_{p}}(G\otimes\mathbb{F}_{p})$.
  • We let $G^{n}$ be the image of the map $G\xrightarrow{\times n}G$ and $G_{n}=\ker(G\xrightarrow{\times n}G)$. We also define $G^{n\ast}\subseteq G$ as the coset
    $\Big{\{}\gamma\mid n\mathfrak{q}(\mu)+(\mu,\gamma)\equiv 0\mod\mathbb{Z},~{}\forall\mu\in G_{n}\Big{\}}.$ (1.2)
  • Let $\mathbb{C}[G^{(d)}]$ be the group algebra of $G^{(d)}=G\times G\ldots\times G$. There is a standard Hermitian inner product $\langle$-,$\rangle$ - , - ⟩ on $\mathbb{C}[G^{(d)}]$ given by
@@
  • Let $(M,\left<-,-\right>)$ be an even lattice. We can identify the dual lattice $M^{\vee}$ as the sublattice
@@
    of $M\otimes\mathbb{Q}$. The discriminant group $G_{M}:=M^{\vee}/M$ is a finite group of order $|\det(M)|$. There is a natural quadratic function $\mathfrak{q}_{M}:M^{\vee}/M\to\mathbb{Q}/\mathbb{Z}$ given by
    $\mathfrak{q}_{M}(v)=\frac{1}{2}\left<v,v\right>\mod\mathbb{Z},$ (1.5)
    for $v\in M^{\vee}$.
  • Given a discriminant form $G$, we use $\mathbf{Gen}_{b^{+},b^{-}}(G)$ to denote the genus of even lattices of signature $(b^{+},b^{-})$ with discriminant form $G$. For a lattice $M\in\mathbf{Gen}_{b^{+},b^{-}}(G)$, the signature of the discriminant form $M^{\vee}/M$ is defined as $\mathrm{sign}(M):=b^{+}-b^{-}\mod 8\in\mathbb{Z}/8\mathbb{Z}$. For all $(b^{+},b^{-})$ such that $\mathbf{Gen}_{b^{+},b^{-}}(G)\neq\varnothing$ and for all $M\in\mathbf{Gen}_{b^{+},b^{-}}(G)$, this quantity is the same. Thus we also use $\mathrm{sign}(G)$ to denote it.
  • We use $\mathbf{e}(x)$ to denote the function $e^{2\pi ix}$, and we use $\sqrt{z}:=z^{\frac{1}{2}}$ to denote the principal branch of the square root function on $\mathbb{C}$.
2 Vector-valued Siegel modular forms
2.1 Weil representation
Let $\mathbb{H}_{d}$ be the Siegel upper half-space. The metaplectic double cover ${\rm Mp}_{2d}(\mathbb{Z})$ of ${\mathrm{Sp}}_{2d}(\mathbb{Z})$ consists of the elements $\widetilde{g}=\left(g,\phi_{g}(\tau)\right)$, where for $g=\left(\begin{array}[]{cc}A&B\\ C&D\end{array}\right)\in{\mathrm{Sp}}_{2d}(\mathbb{Z})$, $\phi_{g}(\tau)$ is a square root of the function $\det(C\tau+D)$. It is well-known that ${\rm Mp}_{2d}(\mathbb{Z})$ is generated by
$J_{d}=\Big{(}\left(\begin{array}[]{cc}0&-\rm{I}_{d}\\ \rm{I}_{d}&0\end{array}\right),\sqrt{\det(\tau)}\Big{)},~{}~{}~{}n(B)=\Big{(}\left(\begin{array}[]{cc}\rm{I}_{d}&B\\ 0&\rm{I}_{d}\end{array}\right),1\Big{)},$

### v2

# The Picard group of the Baily–Borel compactification of the moduli space of quasi-polarized K3 surfaces and generalizations

@@
Abstract.
In this paper, we investigate the Picard group of the Baily–Borel compactification of orthogonal Shimura varieties. As a key result, we determine the Picard group of the Baily–Borel compactification of the moduli space of quasi-polarized K3 surfaces, proving that it is isomorphic to $\mathbb{Z}$. Notably, this contrasts with the moduli space of smooth curves, where the Picard group exhibits a more complex structure after natural compactification.
Our result follows from a general theorem for orthogonal Shimura varieties: for even lattices $M$ of signature $(2,n)$ with $n>8$ satisfying specific arithmetic conditions (e.g. K3 type or $p$-elementary), the rational Picard group of $\overline{\operatorname{Sh}}_{\Gamma}(M)$ with $\Gamma$ containing the stable orthogonal group is $1$-dimensional. The core of our proof lies in constructing an arithmetic obstruction space that governs the extension of Heegner divisors to the generic points of the boundary in orthogonal Shimura varieties. We further establish a connection between this obstruction space and the space of theta series, demonstrating that the obstruction space is maximal under our conditions.
2020 Mathematics Subject Classification:
14J28,14J15,11F27,11F37 (primary), 14G35 (secondary)
Zhiyuan Li is supported by NSFC grant (12121001, 12171090 and 12425105) and Shanghai Pilot Program for Basic Research (No. 21TQ00). Zhiyuan is also a member of LMNS. Manuel K.-H. Müller acknowledges support by Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Centre TRR 326 Geometry and Arithmetic of Uniformized Structures, project number 444845124 as well as support by the LOEWE professorship in Algebra, project number LOEWE/4b//519/05/01.002(0004)/87. Zelin Ye is supported by the NKRD Program of China (No. 2020YFA0713200).
1. Introduction
1.1. Picard group of moduli space of quasi-polarized K3 surfaces
The study of the Picard group of moduli spaces stands as a vital and inspiring challenge in moduli theory, rooted in the groundbreaking contributions of Mumford [Mum65] in 1965. Let ${\mathscr{M}}_{g}$ be the moduli space of smooth curves of genus $g\geq 3$ and let ¯$\overline{{\mathscr{M}}}_{g}$ be its Deligne–Mumford compactification. Harer [Har83] and Arbarello–Cornalba [AC87] showed that
@@ 1.1. Picard group of moduli space of quasi-polarized K3 surfaces
For K3 surfaces, let ${\mathscr{F}}_{g}$ be the (coarse) moduli space of quasi-polarized K3 surfaces of genus $g\geq 2$. By the global Torelli theorem, $\mathscr{F}_{g}$ is realized as an orthogonal Shimura variety, whose rich geometric structure has been extensively studied. Recent breakthroughs [BLMM17, DFV24] established that its Picard group ${\rm Pic}(\mathscr{F}_{g})\cong\mathbb{Z}^{r_{g}}$ is freely generated by linear combinations of Noether–Lefschetz divisors, with rank given by the intricate formula:
$r_{g}=\frac{31g+24}{24}$ $-\frac{1}{4}\left(\frac{2g-2}{2g-3}\right)\left(\frac{g}{2}\right)-\frac{1}{6}\left(\frac{g-1}{4g-5}\right)$
[…]
where $\left(\frac{\cdot}{\cdot}\right)$ denotes the Jacobi symbol. As a locally symmetric variety, ${\mathscr{F}}_{g}$ has a Baily–Borel compactification by the works of Satake, Baily, and Borel, denoted as ¯$\overline{{\mathscr{F}}}_{g}$. A natural problem is to figure out the Picard group of ¯$\overline{{\mathscr{F}}}_{g}$. Since ¯$\overline{{\mathscr{F}}}_{g}$ is a highly singular variety (with canonical singularities), one expects that the Picard rank of ¯$\overline{{\mathscr{F}}}_{g}$ is relatively small. The main result of this paper establishes that the Picard rank of ¯$\overline{\mathscr{F}}_{g}$ is surprisingly minimal—in fact, the smallest possible.
Theorem 1.1.
${\rm Pic}(\overline{{\mathscr{F}}}_{g})\cong\mathbb{Z}$, which is spanned by some multiple of the extended Hodge line bundle $\overline{\lambda}$.
This demonstrates that despite the complexity of ${\rm Pic}({\mathscr{F}}_{g})$, the Baily-Borel compactification exhibits extreme rigidity in its theory of Cartier divisors.
1.2. Picard group of orthogonal Shimura varieties
Our approach establishes a general method for computing Picard groups of Baily–Borel compactifications of orthogonal Shimura varieties. Let $M$ be a lattice of rank $\geq 5$ and ${\rm Sh}_{\Gamma}(M):=\Gamma\backslash\mathbf{D}$ the associated orthogonal Shimura variety, where $\mathbf{D}$ is the Hermitian symmetric domain attached to $M$ and $\Gamma\leq\mathrm{O}(M)$ is a congruence subgroup. Denote by $\overline{{\rm Sh}}_{\Gamma}(M)$ or $(\Gamma\backslash\mathbf{D})^{\rm BB}$ its Baily–Borel compactification. In most situations, we may assume $\Gamma$ is contained in the stable orthogonal group $\widetilde{\mathrm{O}}(M)$.
According to [BLMM17, Thm. 1.5], the rational Picard group $\operatorname{Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))$ is spanned by irreducible components of Heegner divisors on ${\rm Sh}_{\Gamma}(M)$. We define the Heegner subspace $\operatorname{Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))^{\mathrm{Heegner}}$ as the $\mathbb{Q}$-subspace generated by these divisors, and introduce the subgroup
$\operatorname{Pic}(\overline{{\rm Sh}}_{\Gamma}(M))^{\mathrm{Heegner}}\subseteq\operatorname{Pic}(\overline{{\rm Sh}}_{\Gamma}(M))$
consisting of Cartier divisors whose restrictions to $X$ lie in $\operatorname{Pic}_{\mathbb{Q}}(X)^{\mathrm{Heegner}}$.
In this paper, we investigate the problem when elements of $\operatorname{Pic}_{\mathbb{Q}}(\overline{{\rm Sh}}_{\Gamma}(M))^{\mathrm{Heegner}}$ extend to $\overline{{\rm Sh}}_{\Gamma}(M)$ and explicitly compute $\operatorname{Pic}(\overline{{\rm Sh}}_{\Gamma}(M))^{\mathrm{Heegner}}$. Building on Bruinier–Freitag’s analysis of Heegner divisor extensions [BF01], we use theta series and their admissible isotropic lifts to construct an obstruction subspace within the space of cuspidal forms. Under some mild arithmetic conditions on $M$, we prove that this subspace is maximal.
Our main result is encapsulated in the following theorem:
Theorem 1.2 (Cor. 6.7 and 6.10).
Let $M$ be an even lattice of signature $(2,n)$ with $n>8$ satisfying
  • $M$ splits two hyperbolic planes
  • $M\otimes\mathbb{Z}_{p}$ splits three hyperbolic planes for all prime $p$.
For any congruence subgroup $\Gamma$ of $\widetilde{\mathrm{O}}(M)$, we have
${\rm Pic}_{\mathbb{Q}}\left(\overline{{\rm Sh}}_{\Gamma}(M)\right)^{\rm Heegner}\cong\mathbb{Q},$
is spanned by some multiple of the extended Hodge line bundle $\overline{\lambda}$. Moreover, if $\Gamma=\widetilde{\mathrm{O}}(M)$, then
${\rm Pic}\left(\overline{{\rm Sh}}_{\Gamma}(M)\right)^{\rm Heegner}={\rm Pic}\left(\overline{{\rm Sh}}_{\Gamma}(M)\right)\cong\mathbb{Z}.$
Theorem 1.1 is then a special case of Theorem 1.2 applied to the lattice
$\Lambda_{g}\coloneqq\langle 2-2g\rangle\oplus E_{8}(-1)^{\oplus 2}\oplus U^{\oplus 2}.$
@@ 1.2. Picard group of orthogonal Shimura varieties
When $\Gamma$ is neat, the Hodge bundle $\lambda$ extends as a Cartier divisor over $\overline{{\rm Sh}}_{\Gamma}(M)$ (cf. [Pera, Theorem 5.2.11]).
Remark 1.4 (Conditions on $M$).
Let us give some comments on the conditions given in Theorem 1.2. The local and global conditions imposed on $M$ play distinct but crucial roles in our analysis:
  • (Global condition) The condition that $M$ splits two hyperbolic planes guarantees
    1. (a)
      the group ${\rm Pic}({\rm Sh}_{\Gamma}(M))^{\rm Heegner}$ admits a simple description in terms of modular forms;
    2. (b)
      when $\Gamma=\widetilde{\mathrm{O}}(M)$, ${\rm Pic}({\rm Sh}_{\Gamma}(M))={\rm Pic}({\rm Sh}_{\Gamma}(M))^{\rm Heegner}$ is generated by Heegner divisors.
  • (Local condition) Consider $M=L\oplus U^{\oplus 2}$ with $L$ negative-definite. The requirement that $L\otimes\mathbb{Z}_{p}$ splits a hyperbolic plane guarantees surjectivity of certain theta liftings, which generally fails without this condition (see Remark 4.7).
  • For $p$-elementary lattices $M$ which only split a hyperpbolic plane, an analogous result holds when $\Gamma=\mathrm{O}(M)$. The full statement and proof are provided in Theorem 6.4.
1.3. Moduli spaces as Baily–Borel compactifications
A central motivation for our study of the Picard group of Baily–Borel compactifications lies in the ubiquity of these compactifications in moduli theory. Many moduli spaces in algebraic geometry admit natural Baily–Borel compactifications or are themselves isomorphic to such compactifications. This is particularly true for moduli spaces of higher-dimensional varieties with special holonomy, such as hyper-Kähler varieties and some bounded Calabi–Yau surface pairs (cf. [Do96, La10, BL24]).
Recent breakthroughs have established explicit identifications between Baily–Borel compactifications and moduli spaces of boundary polarized Calabi–Yau pairs. In [BL24, Thm. 1.1], Blum–Liu construct a projective asymptotically good moduli space
$\mathscr{M}^{\mathrm{CY}}=\left\{(X,D)~{}~{}\middle|\begin{array}[]{l}\text{(1) }\dim X=2,\,X\text{ projective slc}\\ \text{(2) }K_{X}+D\sim_{\mathbb{Q}}0\\ \text{(3) }D\text{ ample }\mathbb{Q}\text{-divisor}\end{array}\right\}/\sim$
parametrizing $S$-equivalence classes of boundary polarized Calabi–Yau surface pairs. This moduli space (or its seminormalization) serves as a wall-crossing space between K-stability and KSBA compactifications and carries a CM line bundle.
A fundamental example comes from K3 surfaces with non-symplectic involutions. For a K3 surface $S$ with non-symplectic involution $\sigma$ of type $\rho$, the quotient pair
$\left(S/\sigma,\tfrac{1}{2}C\right)$
where $C\subset S$ is the fixed locus of $\sigma$, forms a boundary polarized CY pair. Let ${\mathscr{F}}_{\rho}$ be the coarse moduli space of such pairs, which has been extensively studied in the past few years (see [AE22, AET23, AEH24, AEGS2025, BL24]). There are various modular compactifications of ${\mathscr{F}}_{\rho}$ which fit into a wall-crossing diagram:
$\mathscr{F}^{\mathrm{K}}_{\rho}$$\mathscr{F}^{\mathrm{CY}}_{\rho}$$\mathscr{F}^{\mathrm{KSBA}}_{\rho}$$(\Gamma_{\rho}\backslash\mathbf{D}_{\rho})^{\mathrm{BB}}$-moduliKSBA[BL24, Thm. 1.3]
Here $\mathscr{F}^{\mathrm{CY}}_{\rho}$ is the seminormalization of the closure of $\mathscr{F}_{\rho}$ in $\mathscr{M}^{\mathrm{CY}}$, $\mathbf{D}_{\rho}$ is the type IV Hermitian symmetric domain associated to a lattice $\Lambda_{\rho}$, and $\Gamma_{\rho}=\mathrm{O}(\Lambda_{\rho})$ the orthogonal group. The birational morphism
$(\Gamma_{\rho}\backslash\mathbf{D}_{\rho})^{\rm BB}\to{\mathscr{F}}_{\rho}^{\rm CY}$
is the normalization map. As a by-product of Theorem 6.4, we obtain the following.
Corollary 1.5 (Picard group of BCY moduli spaces).
The rational Picard group of the normalization of $\mathscr{F}^{\mathrm{CY}}_{\rho}$ is spanned by the CM line bundle provided that $\dim\mathscr{F}_{\rho}>8$ and $\mathscr{F}_{\rho}$ is not the 11-dimensional case, in which the fixed locus of $\sigma$ is a smooth rational curve.
The lattice $\Lambda_{\rho}$ is isomorphic to the coinvariant lattice of $\mathrm{H}^{2}(S,\mathbb{Z})$ under the antisymplectic involution, which is a $2$-elementary even lattice. We refer to [Ni79, 3.6.2] and [AE22, 2A] for a complete classification of all $\Lambda_{\rho}$. Crucially, $\Lambda_{\rho}$ need not split two hyperbolic planes—a subtlety that necessitates our approach to Corollary 1.5. This result follows from a dedicated analysis of obstruction spaces and the span of Heegner divisors, employing the study of admissible isotropic lifts of theta series.
@@ 1.3. Moduli spaces as Baily–Borel compactifications
By [Nikulin1983Factorgroups] and see also [AE22], there exist 75 families of K3 surfaces with non-symplectic involutions, of which 38 satisfy $\dim\mathscr{F}_{\rho}>8$. The exceptional case in Corollary 1.5 corresponds to the 11-dimensional family with lattice $\Lambda_{\rho}\cong A_{1}\oplus U(2)\oplus E_{8}(-2)$. Since our general framework does not extend to this case, we leave it open for further study, which may require alternative or more specialized methods.
1.4. The proof strategy
Our main theorem (Theorem 1.1) amounts to characterizing precisely when Heegner divisors extend to the Baily–Borel boundary: this occurs if and only if the divisor class is proportional to the Hodge bundle $\lambda$. The proof synthesizes modular forms and Shimura geometry through three key innovations:
  1. (1)
    (Obstructions theory) A local-to-global criterion which combines Borcherds products [Bo98] with Bruinier–Freitag’s boundary analysis [BF01] establishes the following principle between geometry and arithmetic:
    Geometry Arithmetic
[…]
    Dual of
     Almost Cusp Forms
    Obstruction to Extending
    Heegner divisors
[…]
    (Admissible Isotropic Lifts of)
    theta series
    Roughly speaking, ${\rm Pic}_{\mathbb{Q}}({\rm Sh}_{\Gamma}(M))$ can be identified with the dual of almost cusp forms, while the obstruction space contains the space of theta series and their admissible isotropic lifts, which is a subspace of the space of cusp forms.
  2. (2)
    (Basis problem for modular forms) We study whether the space of cusp forms is generated by theta series or their special isotropic lifts. This is called the basis problem, which is a classical problem in the theory of modular forms and goes back to Eichler. We resolve the basis problem in the vector-valued case for half-integral weight forms (Theorem 4.5), circumventing even signature restrictions in [Ma24].
  3. (3)
    (Torsion-freeness) By extending the work in [DFV24], we establish the torsion-freeness of ${\rm Pic}({\rm Sh}_{\Gamma}(M))$, also simplifying the earlier arguments in [DFV24].
1.5. Organization of the paper
The paper is organized as follows. Section 2 establishes foundations of vector-valued Siegel modular forms, including theta series constructions and Siegel–Eisenstein series, developing essential tools for subsequent sections.
Sections 3 and 4 form the technical core, where we prove surjectivity of theta liftings for odd-rank cases by adapting [Ma24]’s methods through Bruinier–Stein and Shimura Hecke operators. Some computational details are provided in the Appendix, and geometrically-oriented readers may proceed directly to Section 5.
In Sections 5 and 6, we establish the arithmetic obstruction subspace of extending Heegner divisors by
    Extension of Eichler’s criteria to some lattices splitting only one hyperbolic plane, this allows us to generalize some results in [BLMM17];
    Classifying boundary components of Baily–Borel compactifications and introducing so-called admissible isotropic planes, which will be used to construct the obstruction subspace;
    Using Bruinier–Freitag’s local Bocherds product and admissible isotropic lifting to obtain an obstruction subspace for extending Heegner divisors in terms of various theta series.
The main theorem is proved in Section 6, synthesizing these technical components to resolve the Picard group computation, with applications to moduli spaces of Calabi–Yau pairs.
@@ 1.5. Organization of the paper
  • For any normal variety $X$, denote by ${\rm Cl}_{\mathbb{Q}}(X)$ its $\mathbb{Q}$-Weil divisor class group.
@@
    $(\gamma,\delta)\coloneqq\mathfrak{q}(\gamma+\delta)-\mathfrak{q}(\gamma)-\mathfrak{q}(\delta)$
    is a non-degenerate symmetric bilinear form. More generally, for any two elements $\mathbf{u}=(u_{1},...,u_{d}),\mathbf{v}=(v_{1},...,v_{d})\in G^{(d)}=G\times G\times\cdots\times G$, we define $(\mathbf{u},\mathbf{v})\coloneqq\begin{pmatrix}(u_{i},v_{j})_{i,j}\end{pmatrix}\in{\rm Sym}_{d}(\mathbb{Q}/\mathbb{Z})$. The level of $G$ is
    $\mathfrak{n}(G):=\min\{m\in\mathbb{Z}_{>0}\mid~{}m\mathfrak{q}(x)=0\hbox{ for all }x\in G\}.$
    The non-degeneracy of (,)(-,-)( - , - ) implies $\mathfrak{n}(G)\cdot\gamma=0$ for all $\gamma\in G$. The $p$-rank of $G$ is $\mathrm{rank}_{p}(G)\coloneqq\dim_{\mathbb{F}_{p}}(G\otimes\mathbb{F}_{p})$.
  • If $G$ is 2-torsion, one can follow [AE22] to define the coparity of $G$ by $\delta_{G}=0$ if every $2\mathfrak{q}(x)=0\mod\mathbb{Z}$ for all $x\in G$, otherwise $\delta_{G}=1$. There is a unique element $\alpha\in G$ such that
    $2\mathfrak{q}(\gamma)\equiv(\gamma,\alpha)\mod\mathbb{Z}$
    for any $\gamma\in G$, called the characteristic element of $G$. Note that $G$ contains a nonzero, isotropic, characteristic element $\alpha$ if and only if $\delta_{G}=1$ and has signature $0\bmod 4$. In this case, the coparity of $\alpha^{\bot}/\langle\alpha\rangle$ is 00.
  • Define:
    $G^{n}$ $\coloneqq\mathrm{image}(G\xrightarrow{\times n}G),$
    $G_{n}$ $\coloneqq\ker(G\xrightarrow{\times n}G),$
    $G^{n*}$$\coloneqq\left\{\gamma\in G\mid n\mathfrak{q}(\mu)+(\mu,\gamma)\equiv 0\pmod{\mathbb{Z}},\,\forall\mu\in G_{n}\right\}.$
  • Let $\mathbb{C}[G^{(d)}]$ be the group algebra of $G^{(d)}$. There is a standard Hermitian inner product $\langle$-,$\rangle$ - , - ⟩ on $\mathbb{C}[G^{(d)}]$ given by
@@
  • Let $(M,\langle\cdot,\cdot\rangle)$ be an even lattice. We can identify the dual lattice $M^{\vee}$ as the sublattice
@@
    of $M\otimes\mathbb{Q}$. The discriminant group $G_{M}\coloneqq M^{\vee}/M$ satisfies $|G_{M}|=|\det M|$. The quadratic form $\mathfrak{q}_{M}:G_{M}\to\mathbb{Q}/\mathbb{Z}$ is given by
    $\mathfrak{q}_{M}(v)\equiv\tfrac{1}{2}\langle v,v\rangle\mod\mathbb{Z},\quad v\in M^{\vee}.$
    Let $\ell(G_{M})$ denote the minimal number of generators of $G_{M}$.
  • For a discriminant form $G$, let $\mathbf{Gen}_{b^{+},b^{-}}(G)$ denote the genus of even lattices with signature $(b^{+},b^{-})$ and discriminant form $G$. The signature of $G$ is defined as
    $\mathrm{sign}(G)\coloneqq b^{+}-b^{-}\pmod{8}\in\mathbb{Z}/8\mathbb{Z},$
    for $(b^{+},b^{-})$ such that the genus $\mathbf{Gen}_{b^{+},b^{-}}(G)$ is non-empty and the signature is constant over all genera with discriminant form $G$.
  • Define $\mathbf{e}(x)\coloneqq e^{2\pi ix}$ and $\sqrt{z}\coloneqq z^{1/2}$ (principal branch on $\mathbb{C}$).
2. Vector-valued Siegel modular forms
2.1. Weil Representation
Let $\mathbb{H}_{d}$ denote the Siegel upper half-space of degree $d$. The metaplectic double cover ${\rm Mp}_{2d}(\mathbb{Z})$ of ${\mathrm{Sp}}_{2d}(\mathbb{Z})$ consists of elements $\widetilde{g}=\bigl{(}g,\phi_{g}(\tau)\bigr{)}$, where for $g=\begin{pmatrix}A&B\\ C&D\end{pmatrix}\in{\mathrm{Sp}}_{2d}(\mathbb{Z})$, the function $\phi_{g}(\tau)$ is holomorphic on $\mathbb{H}_{d}$ and satisfies $\phi_{g}(\tau)^{2}=\det(C\tau+D)$. The group ${\rm Mp}_{2d}(\mathbb{Z})$ is generated by
$J_{d}=\Bigl{(}\left(\begin{array}[]{cc}0&-\mathrm{I}_{d}\\ \mathrm{I}_{d}&0\end{array}\right),\sqrt{\det\tau}\Bigr{)}\quad\text{and}\quad n(B)=\Bigl{(}\left(\begin{array}[]{cc}\mathrm{I}_{d}&B\\ 0&\mathrm{I}_{d}\end{array}\right),1\Bigr{)},$